OpenStudy (loser66):

\[f(x) =ax_1^2 +2x_1x_2+2x_1x_3\] under which condition of a, f(x) >0? Please

OpenStudy (kinggeorge):

What are \(x_1,x_2,x_3\)?

OpenStudy (loser66):

I am sorry, I edited my question :) I missed 2 before the last term

OpenStudy (loser66):

@KingGeorge actually the problem comes from positive definite matrix. I "translate" it into the simplest form :)

OpenStudy (loser66):

\[\left[\begin{matrix}\alpha&1&1\\1&0&0\\1&0&0\end{matrix}\right]\] question: under which condition of \(\alpha\) the matrix is positive matrix?

OpenStudy (loser66):

and this is translated to \[\alpha x_1^2+2x_1x_2+2x_1x_3\]

OpenStudy (loser66):

if the expression >0 for all x1, x2, x3, then the matrix is positive definite.

ganeshie8 (ganeshie8):

\(\large x_i\) are positive integers ?

OpenStudy (loser66):

or another way to put it in logic is \(x^tAx>0\)

OpenStudy (loser66):

x is an arbitrary vector on vector space

OpenStudy (loser66):

can we put it in logic of x1>0 and the (ax1+2x2 +2x3) > 0 simultaneously to argue for it x1<0 and the ( )<0 x1=0 --> no solution for a is it ok?

ganeshie8 (ganeshie8):

looks okay to me, you want to make \(\large a\) depend on ur vector is it...

OpenStudy (kinggeorge):

We don't want to make \(\alpha\) dependent on the vector.

OpenStudy (kinggeorge):

It seems to me, that there is no \(\alpha\) that will make everything work nicely. For example, if \(\alpha>0\) let \(x_1=1/\sqrt{\alpha}\), \(x_2=0\), and \(x_3=-\sqrt\alpha\). Then\[\alpha x_1^2+2x_1x_2+2x_1x_3=1+0-2=-1\]

OpenStudy (loser66):

A is a Hermitian matrix --> det A1 = alpha > o when alpha>0 det A2 = -1 we can stop here because det A2 <0 for all alpha --> with any value of alpha, A cannot be positive.

OpenStudy (kinggeorge):

If \(\alpha<0\) instead choose \(x_1=1/\sqrt{-\alpha}\), \(x_2=0\), and \(x_3=-\sqrt{-\alpha}\). If \(\alpha=0\), simply choose \(x_1=-1\), and \(x_2=x_3=1\).

OpenStudy (loser66):